OpenStudy (anonymous):

.Choose the equation of the line passing through the point (4, -2) and perpendicular to y = 1/3x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (anonymous):

Firstly find the slope of perpendicular line: y = 1/3x - 7 So coefficient of x will give you slope : 1/3 So the slope of other line will be negative reciprocal of 1/3 that is (-3) So use: \[\large y - y_1 = m(x-x_1)\] So: \[y+2 = -3(x-4)\] So: y = -3x + 12 - 2 y = -3x + 10